OpenStudy (anonymous):

Find the value of y so that the line passing through (9, 0) and (3, y) has a slope of QUICK MEDAL

OpenStudy (alexandervonhumboldt2):

\[\frac{ y_1-y_2 }{ x_1-x_2 }=slope\]

OpenStudy (alexandervonhumboldt2):

just substitute the given values and then solve:

OpenStudy (alexandervonhumboldt2):

what is the value of slope? i dont see it.

OpenStudy (alexandervonhumboldt2):

\[\frac{ 0-y }{ 9-3 }=slope\] slope if given, so just solve

OpenStudy (anonymous):

\[\frac{ 0-y }{ 9-3 } = \frac{ -y }{ 6 }\]

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

what is the value of slope? you haven't written it: your questions says: Find the value of y so that the line passing through (9, 0) and (3, y) has a slope of QUICK MEDAL but what is the vlue of slope?

OpenStudy (anonymous):

1/2

OpenStudy (alexandervonhumboldt2):

ok thus we have \[\frac{ -y }{ 6 }=\frac{ 1 }{ 2 }\] how solve this equation

OpenStudy (anonymous):

i meant -1/2

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (alexandervonhumboldt2):

\[\frac{ -y }{ 6 }=-\frac{ 1 }{ 2 }\]

OpenStudy (alexandervonhumboldt2):

solve this

OpenStudy (anonymous):

i dont know how

OpenStudy (alexandervonhumboldt2):

you have to multply both sides by 6: \[\frac{ -y }{ 6 }*6=-\frac{ 1 }{ 2 }*6\] \[-y=-3\] thus y=3

OpenStudy (anonymous):

ooohhhhhhhhhhh okay i see, wow! thanks